OpenStudy (annie12m):

Medal: 18 = – 2m + 6 – m

Mehek (mehek14):

combine like terms -2m - m = ?

OpenStudy (annie12m):

-3m

Mehek (mehek14):

oke so \(18 = -3m + 6\) move 6 to the other side to move it, you would subtract 18 - 6 = ?

OpenStudy (annie12m):

12

Mehek (mehek14):

\(12 = -3m\) divide by -3 \(\dfrac{12}{-3}=?\)

OpenStudy (annie12m):

-4

Mehek (mehek14):

m = -4

OpenStudy (annie12m):

ty!

Mehek (mehek14):

yw :)